# Adaptive Designs in Fast-Track Registration Processes

### v1

@@ Abstract
Fast-track procedures play an important role in the context of conditional registration of health products, such as conditional approval processes and listing processes for digital health applications. Fast-track procedures offer the potential for earlier patient access to innovative products. They involve two registration steps. The applicants can apply first for conditional registration. A successful conditional registration provides a limited funding or approval period and time to prepare the application for permanent registration, which is the second step of the registration process. For conditional registration, typically, products have to fulfil only a part of the requirements necessary for permanent registration. There is interest in valid and efficient study designs for fast-track procedures. This will be addressed in this paper. A motivating example of the paper is the German fast-track registration process of digital health applications (DiGA) for reimbursement by statutory health insurances. The main focus of the paper is the systematic investigation of the utility of adaptive designs in the context of fast-track registration processes. The paper also covers a careful discussion of the different requirements found in the guidelines and their consequences. We demonstrate that the use of adaptive designs in the context of fast-track processes like the DiGA registration process is, in most cases, much more efficient than the current standard of two separate studies. The results presented in this paper are based on numerical calculations supported by mathematical arguments.
Keywords: Confirmatory adaptive design, efficient study design, digital health application (DiHA), digital therapeutics (DTx), dGA, conditional approval, conditional registration
@@ 1 Introduction
The interest and establishment of the so-called fast-track procedures for the registration of health products and medical devices is currently growing worldwide. Examples are conditional marketing authorisation of the EMA (see [1] and [2]), accelerated approval of the FDA (see [3] and [4]), listing processes for digital health applications and other accelerated approval pathways summarized in [5]. The concept of fast-track procedures is as follows: The applicants have the opportunity to first apply for a conditional registration that provides a limited funding period and time to prepare the application for permanent registration. For conditional registration, typically products have to fulfil only a part of the requirements necessary for permanent registration and some kind of initial evidence. After conditional registration, applicants have a specific maximum period of time to meet all other requirements, including convincing evidence for permanent registration. The aim of fast-track procedures is that patients can benefit quickly from innovative products, particularly if a product fulfils an unmet medical need. Fast-track procedures also strengthen small companies because they can receive funding earlier. Because not all requirements may be fulfilled and fully convincing evidence may still be lacking after conditional registration, the potential risk and costs of a preregistered product must be lower than the benefit of its early availability. An example of a registration process for digital applications is the fast-track procedure for digital health applications (DiGA) in Germany. If a DiGA successfully completes the assessment of the BfArM, it is listed for reimbursement by statutory health insurance companies. Interest in this procedure is currently growing in more and more countries. Different variants of the German fast-track procedure of the BfArM for digital health applications are currently being established or have been established in other countries like France and Austria (see, for instance, [6] and [7]). Our paper is mainly motivated by the German fast-track procedure for DiGA.
@@
In this paper, we consider different design options for fast-track registration processes, whereby we aim to account for the requirements of a conditional registration as well as the requirements for the permanent registration. This will include a careful discussion of the different requirements found in the guidelines, in particular, the BfArM guideline for the registration of DiGA in Germany [8]. This is of interest to applicants, as [9] indicates. Our focus is on overall valid and efficient designs that provide the possibility for a conditional registration and the flexibility with regard to the decision whether to use this possibility or directly strive for a permanent registration. We consider a design as overall valid if it provides overall type I error rate control, in the sense that the probability for a permanent registration of a DiGA without a positive health care effect is bounded by some level $\alpha$, e.g. for the one-sided level of $\alpha=0.025$. Our considerations will include the current standard of separate studies for conditional and permanent registration as well as two-stage adaptive designs (see, for instance, [10]) that permit the use of the data of the conditional registration study also for the permanent registration. To the best of our knowledge, our paper is the first that systematically investigates the utility of adaptive designs in the context of a fast-track registration process. An important aspect of this assessment will be the overall performance of the whole two-step registration process, such as the overall success probability (power) and, as already mentioned, the type I error rate. In particular, we will identify situations where a sufficient overall success probability implies specific constraints for the overall study design. An example of such a constraint is the requirement of sufficiently large sample sizes in either the second or both steps, i.e., for the conditional and/or permanent registration. In addition, we will identify scenarios for which a two-stage registration process is less efficient than the direct application for an permanent registration without an intermediate conditional registration step. However, for some of these situations, a mixture of an adaptive two-stage and single-stage design provides an efficient alternative if we can decide with the first-stage data whether to follow the fast-track process or apply directly for permanent registration. Such options are possible with current DiGA registration regulations in Germany.
@@
Our article is arranged as follows. Section 2 gives a brief overview of the regulations of the established fast-track procedures. Section 3 provides a general discussion of requirements for conditional and permanent registration. Section 4 discusses study designs under the assumption that an permanent registration is impossible or of no further interest after an unsuccessful conditional registration process. Reasons for this restriction could be a lack of resources without conditional registration or regulatory concerns with regard to an initially unsuccessful registration process. In Section 5, we will consider designs with a possibility of a data-dependent decision between the fast-track procedure and direct permanent registration, which guarantee overall type I error rate control. The paper concludes in Section 6 with a detailed summary and discussion of our results as well as their implications.
@@ 2 Examples of current Fast-Track Registration Processes
We briefly outline the legal basis and regulatory requirements of the fast-track procedures for digital health applications established to date. The "app on prescription" in Germany has a pioneering role and was introduced by the Digital Healthcare Act (Ger. Digitale-Versorgung-Gesetz, DVG), which came into force in 2019. Since then, digital health applications (Ger. Digitale Gesundheitsanwendungen, DiGA) can be prescribed by physicians and psychotherapists and are reimbursed by statutory health insurance companies. The currently reimbursable DiGA are listed in the directory of reimbursable digital health applications (DiGA directory [11]). To be listed in the DiGA directory (also called directory here), the applications must successfully complete the assessment of the Federal Institute for Drugs and Medical Devices (Ger. Bundesinstitut für Arzneimittel und Medizinprodukte, BfArM). The details of the registration process and the requirements for reimbursable DiGA are regulated by the Digital Health Applications Ordinance (Ger. Digitale-Gesundheitsanwendungen-Verordnung, DiGAV). A concrete interpretation of that and more practical details can be found in the DiGA guideline [8] of the BfArM. An old guideline version is also available in English; see [12].
@@
There are two different possible ways of permanently being listed in the directory: following the fast-track procedure or applying for a permanent listing directly. The requirements to permanently be listed in the directory are the demonstration of at least one positive healthcare effect (via statistical significance) for a specific patient group as well as all quality and compliance standards from Sections 3 to 6 of the DiGAV. In order to prove a positive healthcare effect, a comparative study must be conducted and submitted to the BfArM. Furthermore, the study must be quantitative and fulfil known scientific standards. It must also be registered in a public trial registry. Compared to the application for a permanent listing, the requirements for a conditional listing are partly less strict because the study for the proof of a positive healthcare effect must not yet have been conducted or concluded. The quality and compliance standards from Sections 3 to 6 DiGAV must still be met. Instead of the proof of a positive healthcare effect via a significant study, the DiGA applicants must submit a so-called systematic evaluation of data. The systematic evaluation of data must analyse data from the use of the DiGA which we call pilot data. The pilot data must plausibly illustrate that a subsequent comparative study will be successful in proving a positive healthcare effect of the DiGA. According to the DiGA guide [8, pp. 109-110], showing this initial evidence includes a relevant (numeric) positive healthcare effect that must be calculated from the pilot data. Further requirements for conditional listing are specified in the DiGA guide. The pilot data collection is intended to prepare the final study for the proof of the positive healthcare effect needed for permanent registration. For instance, this includes case numbers, measuring instruments, endpoints, and recruitment methods. The applicants have to submit the evaluation concept (including study protocol) for the final study when applying for conditional listing. Normally, the patient populations and endpoints investigated in the systematic evaluation of data and the final study should be the same. However, the selection of the primary endpoint for the final study can be done at the end of the pilot data collection. If an applicant receives conditional registration from the BfArM, the DiGA is listed in the directory for one year. Therefore, it is reimbursable by the statutory health insurance companies for at least one year. The proof of a positive healthcare effect via a comparative study must be submitted to the BfArM within the one year of conditional listing. Otherwise, the DiGA is removed from the directory, and the applicant can apply its DiGA for a further listing after one year at the earliest. In summary, the applicants have, with the fast-track procedure, one year more to conduct the comparative study to prove a positive healthcare effect while already being listed as a reimbursable DiGA.
@@
The "app of prescription" in Germany and the fast-track registration process for DiGA have inspired France to establish an analogue regulatory framework for so called Digital Medical Devices, DMD for short, (FR. dispositifs médicaux numériques (DMN)) that are reimbursable by health insurance companies. In 2023, France adopted the fast-track registration process, which is called PECAN (FR. Processus de Certification des Applications Numériques de Santé) with Decree 2023-232 (see, for instance, [13]).
Many other countries are currently in the process of introducing or in discussion of implementing fast-track procedures for digital health applications, such as Belgium, Austria, Italy, and more (see, for instance, [6], Table 1 in [7], [14] or [15]).
@@ 3 Requirements for conditional registration and consequences
This section addresses the regulatory and strategic requirements for conditional and permanent registration. Firstly, we make a few general conventions and assumptions. Usually, the data collection for the conditional and for the permanent registration are realized through two separate studies (see, for instance, [11]). As an alternative, we propose to use two-stage adaptive designs and will investigate their utility. Thus, in the following, the term "stage" is used to refer to the two studies of a design with separate studies and the two stages of an adaptive design. For the separate stage design, we will also use the phrasing "non-adaptive design".
In the below presented numerical investigations, we assume a health care effect parameter $\vartheta$ with $H_{1}:\vartheta>0$ corresponding to a positive healthcare effect and $H_{0}:\vartheta\leq 0$ to an inefficient product. We will further assume that at both stages, $k=1,2$, we obtain a normally distributed estimate $\hat{\vartheta}_{k}\sim\mathcal{N}(\vartheta\,,\,I_{k}^{-1})$ of the healthcare effect with associated p-values $p_{k}=1-\Phi(\hat{\vartheta}_{1}\sqrt{I_{k}})$, where $\Phi$ is the standard normal distribution function and $I_{k}$ is the (Fisher) information of the data from stage $k$. The stage-wise effect estimates are assumed to be calculated from disjoint cohorts and thereby stochastically independent. We further assume that for a permanent registration statistical significance at one-sided level $\alpha$ is required (for the fast track-procedure and direct permanent registration). In our numerical investigation we assume the commonly required level $\alpha=0.025$. A possible additional requirement for permanent registration in the case of the fast-track procedure could be that the application for conditional registration must have been successful (for whatever reason). We will leave this open for now and discuss the fast-track procedure in this section with and without this additional requirement.
@@ 3.1 Requirements motivated by the German DiGA guide
For the conditional registration, motivated by the German DiGA guide for digital health applications, we assume that there is a minimal clinically relevant effect $\delta_{\text{rel}}>0$, which is strict in the sense that a product with a healthcare effect lower than $\delta_{\text{rel}}$ is considered as not having a relevant positive impact. Accordingly, a natural requirement for the conditional registration is that, at stage one, a point estimator of $\hat{\vartheta}_{1}\geq\delta_{\text{rel}}$ is observed.
@@
Another natural requirement for conditional registration is $p_{1}\leq\alpha_{c}$ for some significance level $\alpha_{c}\in(0,1)$. This bounds the probability for a temporary registration of an inefficient product. There are several reasons why a conditional registration of an inefficient product is unfavourable, also when the final study for a permanent registration fails. One reason are the costs coming with a temporary registration without a relevant healthcare effect. Users may even be exposed to a product with negative healthcare or side effects during the temporary period of registration. In addition, applicants and health insurance companies may suffer a considerable loss of reputation. This requirement is also mentioned in the German DiGA guide, whereby a two-sided significance level of $0.1$ is indicated as a possibility for conditional registration [8, p. 110], which corresponds to the one-sided stage-one level of $\alpha_{c}=0.05$.
@@
We list further possible requirements for conditional registration. In the DiGA guide, it is indicated that for a successful conditional registration, the study plan for the final stage must be submitted. In this context, it must plausibly be illustrated that a subsequent comparative study/stage will be successful in proving that the DiGA has a positive healthcare effect. This requirement is justified by the costs and complications (for provider and patients) form a withdrawal of registration, which also comes with a loss of reputation. We propose the following specification of this requirement: The power calculation for the second stage should be done for a sufficiently high power like $80\%$ (or higher) at the observed first stage healthcare effect or a lower value.
@@
If the application for a permanent registration is impossible after an unsuccessful study for conditional registration, the impact of the requirements $\hat{\vartheta}_{1}\geq\delta_{\text{rel}}$ and $p_{1}\leq\alpha_{c}$ for a conditional registration is substantial and therefore requires a careful investigation. Therefore, we will investigate these regulatory requirements and their impact on strategic requirements of the applicants like power considerations.
@@
A priori, it is not clear which of the above mentioned two requirements is more stringent. It can be calculated that $\hat{\vartheta}_{1}\geq\delta_{\text{rel}}$ is equivalent to $p_{1}\leq\alpha_{\text{rel}}:=1-\Phi(\delta_{\text{rel}}\sqrt{I_{1}})$. Hence, the first requirement is equivalent to the second one with a specific stage-one level. We introduce the following notations to visualise this level in the relative scale:
@@
$t_{\text{rel}}(I_{1}):={I_{1}}/{I_{\text{rel}}}\quad\text{with}\quad I_{\text{rel}}:={\eta_{f}^{2}}/{\delta_{\text{rel}}^{2}}\text{ where }\eta_{f}:=\Phi^{-1}(1-\beta)+\Phi^{-1}(1-\alpha)\leavevmode\nobreak\ .$ (1)
@@
The parameter $\eta_{f}$ defines the noncentrality parameter and $I_{\text{rel}}$ the information needed to achieve a power of $1-\beta$ at $\delta_{\text{rel}}$ with significance level $\alpha$. In our numerical investigations we assume the commonly used $\alpha=0.025$ (one-sided) and $0.81-\beta=$ 0.8, because these values are used in the majority of DiGA studies for permanent registration. This leads to $\eta_{f}=2.8$. The significance level $\alpha_{\text{rel}}$ is shown in Figure 1 (black solid decreasing line) in dependence on $t_{\text{rel}}(I_{1})$. The $t_{\text{rel}}(I_{1})$-axis ends at $1$, because for $t_{\text{rel}}(I_{1})>1$, it does not make sense to decide in favour of the fast-track procedure: A first stage information $I_{1}>I_{\text{rel}}$ ensures a power of at least $1-\beta$ for rejecting $H_{0}$ at level $\alpha$ with a single fixed-size sample test for all $\delta\geq\delta_{\text{rel}}$. In this case, a study for a direct, permanent registration is more efficient than the pilot study to achieve conditional registration only. Thus $I_{\text{rel}}$ yields an kind of universal upper bound for $I_{1}$.
@@
Figure 1: Type I error rate of conditional registration resulting from the requirement $\hat{\vartheta}_{1}\geq\delta_{\text{rel}}$ (black decreasing line) in comparison to levels $\alpha_{c}=0.15,0.1,0.05$ (one-sided) for conditional registration (green, red and blue horizontal lines).
@@
in the pilot study. The levels $\alpha_{c}=0.05,0.1,0.15$ (one-sided) are indicated in Figure 1 by horizontal lines. For the different choices of $\alpha_{c}$, the Type I error rate $\alpha_{f}$ for conditional registration can be read off Figure 1 by using the minimum of the corresponding horizontal and decreasing black line. For example, with the choice $\alpha_{c}=0.05$ indicated in [8, p. 110] and $I_{1}\leq 0.3447\cdot I_{\text{rel}}$, the requirement for a type I error rate control at level $\alpha_{c}$ is more stringent than the requirement of a point estimator above $\delta_{\text{rel}}$.

### v3

@@ Abstract
Fast-track procedures play an important role in the context of conditional registration of medical devices, such as listing processes for digital health applications. They offer the potential for earlier patient access to innovative products and involve two registration steps. The applicants can apply first for conditional registration. A successful conditional registration provides a limited funding or approval period and time to prepare the application for permanent registration (the second registration step). For conditional registration, products have to fulfill only a part of the requirements necessary for permanent registration. There is interest in valid and efficient study designs for fast-track procedures. This will be addressed in this paper. A motivating example is the German fast-track registration process of digital health applications (DiGA) for reimbursement by statutory health insurances. The main focus of the paper is the systematic statistical investigation of the utility of adaptive designs in the context of fast-track registration processes like the DiGA fast-track. We demonstrate that, in most cases, such designs are much more efficient than the current standard of two separate studies. A careful statistical discussion of the registration requirements and their consequences is also included. The results are based on numerical calculations supported by mathematical arguments.
Keywords: Confirmatory adaptive design, efficient study design, digital therapeutics (DTx), conditional approval, conditional registration
@@ 1 Introduction
The interest in the so-called fast-track procedures for the registration of health products and medical devices is currently growing worldwide. Examples are listing processes for digital health applications, conditional marketing authorisation of the EMA [1, 2], accelerated approval of the FDA [3, 4], and other accelerated approval pathways summarized in the work of Gupte et al [5]. The concept of fast-track procedures is as follows: The applicants have the opportunity to first apply for a conditional registration that provides a limited funding period and time to prepare the application for permanent registration. For conditional registration, products have to fulfill less strict requirements than those necessary for permanent registration, e.g., with regard to the evidence. After conditional registration, applicants have a specific maximum period of time to meet all the requirements for permanent registration, including convincing evidence. The aim of fast-track procedures is that patients can benefit quickly from innovative products, particularly if a product fulfills an unmet medical need. Fast-track procedures are currently becoming increasingly important in the field of digital medical devices. There are many countries that have established such procedures for digital medical devices. This includes countries within (like France, Germany, Belgium) and also outside the EU (like Korea and the United Kingdom), as summarized in an OECD Working Paper of Chapman [6]. There are also many countries that are in the discussion of implementing fast-track procedures, see for instance Tarricone et al [7] or [8]. Furthermore, the existing registration procedures have also become a sales market for foreign companies like Czechia, Poland, and the United States [9, 10].
Our paper is mainly motivated by the German fast-track registration procedure for digital health applications (DiGA), as it has been and still is a pioneer for the establishment of fast-track programs for digital medical devices in many other countries [11]. If a DiGA successfully completes the assessment of the BfArM, it is registered for reimbursement by statutory health insurance companies in Germany. The Federal Institute for Drugs and Medical Devices (BfArM) is the German regulatory authority responsible for medicines and digital health applications.
@@
In this paper, we consider different design options for fast-track registration processes, whereby we account for the requirements of a conditional registration as well as the requirements for permanent registration. On the one hand, this provides a guide for applicants on how to design efficient studies for a fast-track registration process. On the other hand, this will include a careful statistical discussion of the different requirements found in the guidelines, in particular, the BfArM guideline for the registration of DiGA in Germany [12] (an old BfArM guideline version is also available in English [13]). The discussion of these requirements is important, as governments and health insurance companies of many countries are increasingly faced with the challenge of deciding which health technologies they should pay for [6], and on the other side, the community of applicants wants fair requirements as it has been indicated in a survey by Ataiy et al [14]. Taking into account the discussed requirements, our focus is on overall valid, efficient designs. We consider a design as overall valid if it provides overall type I error rate control, in the sense that the probability for a permanent registration of a DiGA without a positive effect in the considered endpoint describing the potential health care effect is bounded by some level $\alpha$, e.g. for the one-sided level of $\alpha=0.025$. Our considerations will include the current standard of separate studies for conditional and permanent registration, as well as two-stage adaptive designs (see, for instance, the work of Wassmer and Brannath [15]) that permit the use of the data of the conditional registration study also for the permanent registration. To the best of our knowledge, this is the first paper to systematically investigate the usefulness of adaptive designs in the context of a fast-track registration process for digital health applications. We illustrate that under common scenarios, adaptive designs are much more efficient than the standard approach of two separate studies for conditional and permanent registration. An important aspect of this assessment will be the overall performance of the whole two-step registration process, such as the overall success probability (power to receive permanent registration) and, as already mentioned, the type I error rate. One important result of our requirements discussion is the identification of scenarios for which a two-stage registration process is less efficient than the direct application for a permanent registration without an intermediate conditional registration step. For such situations, we suggest a data-driven choice based on the first-stage data between the fast-track procedure and the direct application for permanent registration. In particular, we will demonstrate that a mixture of an adaptive two-stage and single-stage design can provide an efficient alternative to applying for permanent registration directly. Such options are possible with the current DiGA registration regulations in Germany.
@@
Our article is arranged as follows. Section 2 gives a brief overview of the regulations and the established fast-track procedures of the different countries. Section 3 provides a general discussion of requirements for conditional and permanent registration. Section 4 discusses study designs under the assumption that a permanent registration is impossible or of no further interest after an unsuccessful conditional registration process. Reasons for this restriction could be a lack of resources without conditional registration or regulatory concerns with regard to an initially unsuccessful registration process. In Section 5, we will consider designs with a possibility of a data-dependent decision between the fast-track procedure and direct permanent registration, which guarantee overall type I error rate control. The paper concludes in Section 6 with a detailed summary and discussion of our results as well as their implications.
@@ 2 Examples of current Fast-Track Registration Processes
We briefly outline the legal basis and regulatory requirements of the fast-track procedures for digital medical devices established to date. The “app on prescription” in Germany has a pioneering role and was introduced by the Digital Healthcare Act, which came into force in 2019. Since then, digital health applications (Ger. Digitale Gesundheitsanwendungen, DiGA) can be prescribed by physicians and psychotherapists and are reimbursed by statutory health insurance companies. The currently reimbursable DiGA are listed in the DiGA directory [9]. To be listed in the directory, the applications must successfully complete the assessment of the BfArM. The details of the registration process and the requirements for reimbursable DiGA are regulated by the Digital Health Applications Ordinance (DiGAV). A concrete interpretation of that and more practical details can be found in the DiGA guideline of the BfArM [12]. An old guideline version is also available in English [13].
@@
There are two different possible ways of permanently being listed in the directory: following the fast-track procedure or applying for a permanent listing directly. The requirements to permanently be listed in the directory are the demonstration of at least one positive healthcare effect (via statistical significance) for a specific patient group, as well as all quality and compliance standards from Sections 3 to 6 of the DiGAV. In order to prove a positive healthcare effect, a quantitative comparative study should be conducted and submitted to the BfArM. It must also be registered in a public trial registry. Compared to the application for a permanent listing, the requirements for a conditional listing are partly less strict because the study for the proof of a positive healthcare effect must not yet have been conducted or concluded. Instead, the DiGA applicants must submit a so-called systematic evaluation of data where data from the use of the DiGA must be analyzed. This data, which we call pilot data, must plausibly illustrate that a subsequent comparative study will be successful in proving a positive healthcare effect of the DiGA [12, pp. 109-110]. According to the DiGA guide, showing this initial evidence includes a relevant (numeric) positive healthcare effect that must be calculated from the pilot data [12, p. 110]. Further requirements are specified in the DiGA guide. The pilot data collection is intended to prepare the final study for the proof of the positive healthcare effect needed for permanent registration. For instance, this includes case numbers, measuring instruments, endpoints, and recruitment methods. The applicants have to submit the evaluation concept (including study protocol) for the final study when applying for conditional listing. Ideally, the patient populations and endpoints investigated in the systematic evaluation of data and the final study should be the same. If an applicant receives conditional registration from the BfArM, the DiGA is listed for reimbursement in the directory for one year. The proof of a positive healthcare effect via a comparative study must be submitted to the BfArM within the one year of conditional listing. Otherwise, the DiGA is removed from the directory, and the applicant can apply its DiGA for a further listing after one year at the earliest. In summary, the applicants have, with the fast-track procedure, one year more to conduct the comparative study to prove a positive healthcare effect while already being listed as a reimbursable DiGA.
@@
The “app on prescription” in Germany and the fast-track registration process for DiGA have inspired France to establish an analogue regulatory framework for digital medical devices, DMD for short, (FR. dispositifs médicaux numériques (DMN)) that are reimbursable by health insurance companies. In 2023, France adopted the fast-track registration process, which is called PECAN (FR. Processus de Certification des Applications Numériques de Santé) [16] with Decree 2023-232. Belgium has also established an authorisation pathway for digital health applications (FR. Applications de santé mobiles) with the mHealth pyramid, see [10], via the royal decree [17]. Many other EU countries are currently in the process of introducing or in discussion of implementing fast-track procedures for digital health applications, such as Austria, Italy, and more [7, 8, 18]. There are also ambitions to create a uniform assessment framework for digital interventions within the EU via the ASSESS DHT [19] or EDiHTA [20] project.
Outside the EU, for instance, in Korea, the fast-track pathway Integrated Review and Assessment Programme for Innovative Medical Devices was established in 2022 and 2023, which enables provisional insurance listing based on initial evidence [6]. In the UK, there is the NICE’s Early Value Assessment (EVA) pathway, a fast-track process for (digital) technologies which offers the possibility of a conditional recommendation which can provide a basis for conditional funding [6, 21].
@@ 3 Requirements for conditional registration and consequences
This section addresses the regulatory and strategic requirements for conditional and permanent registration motivated by the German fast-track procedure. Firstly, we make a few general assumptions. Usually, the data collection for the conditional and for the permanent registration is realized through two separate studies [9]. In this paper, we consider the scenario in which the two separate studies have the same design, with the same primary endpoint and the same study population. Later, we propose using two-stage adaptive designs as an alternative to non-adaptive separate study designs. In the following, the term “stage” is used to refer to the two studies of a non-adaptive design and the two stages of an adaptive design.
In the numerical investigations presented below, we assume a health care effect parameter $\vartheta$ with $H_{1}:\vartheta>0$ corresponding to a positive healthcare effect and $H_{0}:\vartheta\leq 0$ to an inefficient product. We will further assume that at both stages, $k=1,2$, we obtain a normally distributed estimate $\hat{\vartheta}_{k}\sim\mathcal{N}(\vartheta\,,\,I_{k}^{-1})$ of the healthcare effect with associated one-sided p-values $p_{k}=1-\Phi(\hat{\vartheta}_{1}\sqrt{I_{k}})$, where $\Phi$ is the standard normal distribution function and $I_{k}$ is the (Fisher) information of the data from stage $k$. Loosely speaking, the information measures how precisely a parameter can be estimated from the data — the larger it is, the more precise the estimate. It is typically proportional to the sample sizes. The stage-wise effect estimates are assumed to be calculated from disjoint cohorts and thereby stochastically independent. We further assume that for a permanent registration, statistical significance at one-sided level $\alpha=0.025$ is required (for the fast-track procedure and direct permanent registration). We assume that applicants must announce before conducting a stage/study, whether it is intended for conditional or permanent registration, in order to avoid inflation of the Type I error rate $\alpha$.
A possible additional requirement for permanent registration, besides statistical significance, in the case of the fast-track procedure, could be that the application for conditional registration must have been successful (for whatever reason). We will leave this open for now and discuss the fast-track procedure in this section with and without this additional requirement.
@@ 3.1 Requirements motivated by the German DiGA guide
For the conditional registration, motivated by the German DiGA guide, we assume that there is a minimal clinically relevant effect $\delta_{\text{rel}}>0$, which is strict in the sense that a product with a healthcare effect lower than $\delta_{\text{rel}}$ is considered as not having a relevant positive impact. Accordingly, a natural requirement for the conditional registration is that, at stage one, a point estimator of $\hat{\vartheta}_{1}\geq\delta_{\text{rel}}$ is observed.
@@
Another natural requirement for conditional registration is $p_{1}\leq\alpha_{c}$ for some significance level $\alpha_{c}\in(0,1)$. This bounds the probability for a temporary registration of an ineffective product. There are several reasons why the conditional registration of an ineffective product is undesirable, even when the final study for a permanent registration fails. One reason are the costs coming with a temporary registration without a relevant healthcare effect. Users may even be exposed to a product with negative healthcare or side effects during the temporary period of registration. In addition, applicants and health insurance companies may suffer a considerable loss of reputation. This requirement is also mentioned in the German DiGA guide, whereby a two-sided significance level of $0.1$ is indicated as a possibility for conditional registration [12], which corresponds to the one-sided stage-one level of $\alpha_{c}=0.05$.
@@
We list further possible requirements for conditional registration. In the DiGA guide, it is indicated that for a successful conditional registration, the study plan for the final stage must be submitted. In this context, it must plausibly be illustrated that a subsequent comparative study/stage will be successful in proving that the DiGA has a positive healthcare effect. This requirement is justified by the costs and complications (for the provider and patients) from a withdrawal of registration, which also comes with a loss of reputation. We propose the following specification of this requirement: The power calculation for the second stage should be done for a sufficiently high power, like $80\%$ (or higher) at the observed first stage healthcare effect, or a lower value.
@@
If the application for a permanent registration is impossible after an unsuccessful study for conditional registration, the impact of the requirements $\hat{\vartheta}_{1}\geq\delta_{\text{rel}}$ and $p_{1}\leq\alpha_{c}$ for a conditional registration is substantial and therefore requires a careful investigation. Therefore, we will investigate these regulatory requirements and their impact on the strategic requirements of the applicants, such as power considerations.
@@
A priori, it is not clear which of the above-mentioned two requirements is more stringent. It can be calculated that $\hat{\vartheta}_{1}\geq\delta_{\text{rel}}$ is equivalent to $p_{1}\leq\alpha_{\text{rel}}:=1-\Phi(\delta_{\text{rel}}\sqrt{I_{1}})$. Hence, the first requirement is equivalent to the second one with a specific stage-one level. We introduce the following notations to visualize this level in the relative scale:
@@
$t_{\text{rel}}(I_{1}):={I_{1}}/{I_{\text{rel}}}\quad\text{with}\quad I_{\text{rel}}:={\eta_{f}^{2}}/{\delta_{\text{rel}}^{2}}\text{ where }\eta_{f}:=\Phi^{-1}(1-\beta)+\Phi^{-1}(1-\alpha)\penalty 10000\ .$ (1)
@@
The parameter $\eta_{f}$ defines the noncentrality parameter and $I_{\text{rel}}$ the information needed to achieve a power of $1-\beta$ at $\delta_{\text{rel}}$ with significance level $\alpha$. In our numerical investigations, we assume the commonly used $\alpha=0.025$ (one-sided) and $1-\beta=0.8$, because these values are used in the majority of DiGA studies for permanent registration. This leads to $\eta_{f}=2.8$. The significance level $\alpha_{\text{rel}}$ is shown in Figure 1 (solid decreasing line) in dependence on $t_{\text{rel}}(I_{1})$. The $t_{\text{rel}}(I_{1})$-axis ends at $1$, because for $t_{\text{rel}}(I_{1})>1$, it does not make sense to decide in favor of the fast-track procedure: A first stage information $I_{1}>I_{\text{rel}}$ ensures a power of at least $1-\beta$ for rejecting $H_{0}$ at level $\alpha$ with a single fixed-size sample test for all $\delta\geq\delta_{\text{rel}}$. In this case, a study for a direct, permanent registration is more efficient than the pilot study to achieve conditional registration only. Thus $I_{\text{rel}}$ can be considered as a universal upper bound for $I_{1}$.
@@
Figure 1: Type I error rate of conditional registration resulting from the requirement $\hat{\vartheta}_{1}\geq\delta_{\text{rel}}$ in comparison to different levels $\alpha_{c}$ (one-sided).
@@
in the pilot study. The levels $\alpha_{c}=0.05,0.1,0.15$ (one-sided) are indicated in Figure 1 by horizontal lines. For the different choices of $\alpha_{c}$, the Type I error rate $\alpha_{f}$ for conditional registration can be read off Figure 1 by using the minimum of the corresponding horizontal and decreasing solid line. For example, with the choice $\alpha_{c}=0.05$ indicated in the DiGA guide [12] and $I_{1}\leq 0.3447\cdot I_{\text{rel}}$, the requirement for a type I error rate control at level $\alpha_{c}$ is more stringent than the requirement of a point estimator above $\delta_{\text{rel}}$.